OpenStudy (anonymous):

13(7)+82 how do I figure that out what do I do with the 7

myininaya (myininaya):

13(7) means multiply

OpenStudy (anonymous):

so multiply 13x7 and add 82

myininaya (myininaya):

yes you just use order of operations

OpenStudy (anonymous):

so I did 13x7=91+82=173

myininaya (myininaya):

well actually you did 13(7)+82=91+82=173 so basically that line that you have is right

OpenStudy (anonymous):

ok thanks a lot